1. On January 1, 2019, Melrose Manufacturing issued a 5-year bond with a face value of $20,000 and a stated interest rate of 6%. The market interest rate is 4% and the bond was issued for $21,780.73. Using the effective-interest method of amortization, how much interest expense will be recorded for the year ended December 31, 2020? A. $871.23 B. $1,200.00 C. $844.40 D. $858.08
Answer #1

Answer : D) $858.08

Interest Expenses = Bond Carrying Value * Market interest rate = 21,708.73*4% =871.33

Cash interest = 20,000*6% = 1200

Premium aomortized = 1200-871.33= 328.67

Bond interest Expenses for 2020 = (21780.73-328.67)*4% = $858.08 (answer)